Circular No. 9130

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

V407 CYGNI The Bureau received reports by Japanese amateurs (K. Nishiyama and

F. Kabashima; T. Kojima; K. Sakaniwa; and A. Tago) of a possible new 7thmag nova in Cygnus on wide-field camera-lens CCD images taken during Mar. 10.8–11.8 UT. Nishiyama and Kabashima suggested that this might be an unusually bright outburst of V407 Cyg (their astrometry yielding α = 21^{h}02^{m}09^{s}.83, δ = +45^{o}46′33′′.0, equinox 2000.0) — not known to have been brighter than mag 13 since its 1936 discovery; it had been classified as a Mira variable with a symbiotic-star spectrum (e.g., Duerbeck 1987, Space Sci. Rev. 45, 1). Visual magnitude estimates by P. Schmeer, Bischmisheim, Germany: Mar. 12.184, 7.9; 17.183, 8.9; 18.191, 9.1. U. Munari et al. report (CBET 2204) that spectra from Mar. 13.09 show the absorption spectrum of the Mira star that is completely overwhelmed by the blue continuum from the outbursting component, a scenario reminiscent of the symbiotic recurrent nova RS Oph; the broad component of the Balmer emission series and He I lines looks closely similar to the spectrum of He/N classical novae. Additional details appear on CBETs 2199, 2205, and 2210.

COMET P/2002 FA_{9} = P/2010 F2 (LINEAR) R. H. McNaught has rediscovered 2002 FA_{9}, an apparently asteroidal

object discovered by the LINEAR survey last year (discovery observation, from MPS 55865, tabulated below), on CCD images taken with the 0.5-m Uppsala Schmidt telescope (McNaught’s first observation from 2010 Mar. 20 is also tabulated below). McNaught found a slightly diffuse 10′′ coma on the 20-s discovery images in mediocre seeing; three stacked 60-s follow-up images show a 16′′ coma extending to the north. The orbit on MPO 30730 requires a correction of ΔT = −0.8 day.

Date UT α_{2000} δ_{2000} Mag. Observer
2002 Mar. 16.31907 11^{h}45^{m}44.^{s}89 −5^{o}23′01′′.8 18.5 LINEAR
2010 Mar. 20.60759 11 40 25.00 −4 38 19.8 18.3 McNaught

The new astrometry, the following orbital elements, and an ephemeris appear on MPEC 2010-F50.

Epoch = 2010 Mar. 25.0 TT

T = 2010 Mar. 21.6348 TT  ω = 333.^{o}7142
e = 0.313683  Ω = 204.4870 }2000.0
q = 2.748317 AU  i = 8.8899

a = 4.004446 AU n^{o} = 0.1229958 P = 8.01 years

2010 March 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green